OpenStudy (anonymous):

If a factory continuously pumps pollutants into the air at the rate of the quotient of the square root of t and 15 tons per day, then the amount dumped after 3 days is

OpenStudy (anonymous):

\[\frac{ \sqrt{t} }{ 15 }\]

OpenStudy (anonymous):

@IrishBoy123 @triciaal

OpenStudy (anonymous):

u need me bro

OpenStudy (anonymous):

umm yeah

OpenStudy (anonymous):

wait was that a question or a statement ?

OpenStudy (anonymous):

it is 45 tons in three days

OpenStudy (anonymous):

sorry its 6.708203932

OpenStudy (anonymous):

Medal plzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (irishboy123):

\(\large \frac{dP}{dt} = \frac{\sqrt{t}}{15} \) \(\large P(3) = \frac{1}{15} \ \int_{t=0}^{3} \sqrt{t} \ dt = \frac{1}{15} | \frac{2}{3}.t^{3/2}|_{t=0}^{3} = \frac{2}{45}. \sqrt{27} =\frac{2 \sqrt{3}}{15} = 0.231\)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

now im confused

OpenStudy (irishboy123):

@watchpencilpaper is this what you were expecting? did you have a method in mind for solving this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i

OpenStudy (anonymous):

did

OpenStudy (anonymous):

have

OpenStudy (anonymous):

a

OpenStudy (anonymous):

method

OpenStudy (anonymous):

no i had no method in mind and it appears you are right @IrishBoy123

OpenStudy (irishboy123):

@watchpencilpaper so how were you going to solve this -- without using calculus?!?!

OpenStudy (irishboy123):

@Napolions your method please!! please share it.

OpenStudy (anonymous):

what do you mean, i intended to say i did not know how to solve it, obviously you need calculus

OpenStudy (irishboy123):

ok @Napolions please, is there a cooler way to do this? do share.

OpenStudy (anonymous):

naw but there was a website that helped me